OpenStudy (anonymous):

Solve. (x – 5)^2 = 4. x = 3 and x = 11 x = 9 and x = 11 x = 9 and x = 1 x = 7 and x = 3

OpenStudy (anonymous):

x=3,7

OpenStudy (anonymous):

x^2-10x+25=4 x^2-10x+21=0 (x-3)(x-7) x=3,7